# Stochastic Optimization with Optimal Importance Sampling

### v2

@@ 5.2 Identification of Active and Inactive Constraints
We will now show that the procedure (19) identifies the active constraints. This proof is a straightforward application of Lemma A.2, as explained in what follows. Following a similar reasoning as in [duchi2021asymptotic, Theorem 3], we start by rewriting iteration (19) as
@@ 5.3 Asymptotic Normality
Assumption 5.6(i) is a standard second-order regularity assumption needed to prove the asymptotic normality of SA algorithms [polyak1992acceleration]. In our case, since we are dealing with a joint SA procedure, Assumption 5.6(ii) is a natural symmetric requirement for the $\mu_{n}$ iterates. Assumption 5.6(iii) is a standard restricted strong convexity assumption needed to prove the asymptotic normality of both the SAA and SA procedures [shapiro1989asymptotic, duchi2021asymptotic].
@@
Assumption 5.7 is a common assumption needed to ensure that the noise sequence $\{\xi_{k}\}_{k\in\mathbb{N}}$ satisfies the CLT $\frac{1}{\sqrt{n}}\sum_{k=0}^{n-1}\xi_{k}\overset{d}{\to}\mathcal{N}(0,\Sigma)$, for some $\Sigma\geq 0$. In particular, Assumption 5.7(i) is the standard conditional Lindeberg condition needed to prove martingale CLTs (see [hall2014martingale, Chapter 3]). Moreover, Assumption 5.7(ii) guarantees that $\Sigma$ is a block-diagonal matrix with $\text{Var}_{X^{(\mu^{\star})}\sim\mathbb{P}_{\mu^{\star}}}\left[G_{\mu^{\star}}(\theta^{\star},X^{(\mu^{\star})})\right]$ and $\text{Var}_{X\sim\mathbb{P}}\left[H(\theta^{\star},\mu^{\star},X)\right]$ on the diagonal.
@@ Acknowledgments
Liviu Aolaritei acknowledges support from the Swiss National Science Foundation through the Postdoc.Mobility Fellowship (grant agreement P500PT_222215). Michael I. Jordan was funded in part by the European Union (ERC-2022-SYG-OCEAN-101071601). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.
@@ Appendix A Technical Preliminaries
Lemma A.1 (robbins1971convergence).
@@
Lemma A.2 (duchi2021asymptotic, Lemma 4.2).
@@
Lemma A.3 (dembo2016lecture, Exercise 5.3.35).
@@
Lemma A.4 (hall2014martingale, Corollary 3.1).
@@ A.1 A Generic Asymptotic Normality Result
To keep the paper self-contained, in what follows we recall the generic asymptotic normality result from Section 13 (pages 12-13) in the Supplementary material of [duchi2021asymptotic]. There, Duchi and Ruan generalize Polyak and Juditsky’s results [polyak1992acceleration] on asymptotic normality in averaged stochastic gradient methods restricted to an arbitrary subspace of $\mathbb{R}^{s}$. Before stating the result, we need to introduce some notation, as well as two assumptions.
@@
We would like to highlight that Assumption A.5(ii) is slightly weaker compared to Assumption S.A in Section 13 in the Supplementary material of [duchi2021asymptotic]. Under Assumption 4.1(iii), this weaker version will be enough for us to prove the CLT in Theorem 5.9.
@@
Lemma A.7 (duchi2021asymptotic, Proposition S.1).
@@
We highlight the difference between the asymptotic variance in (47) and the asymptotic variance in [duchi2021asymptotic, Proposition S.1 in Supplementary material], which is $(\text{P}H\text{P})^{\dagger}\text{P}\Sigma\text{P}(\text{P}H\text{P})^{\dagger}$. The two expressions are equal. This follows from the fact that P is an orthogonal projection matrix, which guarantees that $\text{P}(\text{P}H\text{P})^{\dagger}=(\text{P}H\text{P})^{\dagger}=(\text{P}H\text{P})^{\dagger}\text{P}$.
@@
The proof follows along similar lines as duchi2021asymptotic. For completeness, and since these results are fundamental for proving Theorem 5.2, we provide a full proof for Lemma B.1.
@@
with $\delta_{\Theta}$ the indicator function of the set $\Theta$. Defining $\ell(\theta):=\frac{1}{2}\|\theta\|^{2}+\delta_{\Theta}(\theta)$, we have that the above maximum is precisely the convex conjugate $\ell^{*}(-\sum_{k=0}^{n}\alpha_{k+1}G_{k})$. Now, since $\min_{\theta\in\mathbb{R}^{s}}\{\langle\sum_{k=0}^{n}\alpha_{k+1}G_{k},\theta\rangle+\frac{1}{2}\|\theta\|^{2}+\delta_{\Theta}(\theta)\}$ is the Moreau envelope of $\sum_{k=0}^{n}\alpha_{k+1}G_{k}+\delta_{\Theta}(\theta)$ evaluated at zero, we have that the gradient $\nabla\ell^{*}(-\sum_{k=0}^{n}\alpha_{k+1}G_{k})$ is equal to proximal map evaluated at zero [rockafellar2009variational, Theorem 2.26], i.e.,
@@
The proof follows from [hall2014martingale, Corollary 3.1], using Assumption 5.7 and the Cramér-Wold theorem. Fix $t\in\mathbb{R}^{s+m}$. Then, recalling the shorthand notation $\xi_{k}$ for the noise vector, we have that

### v3

@@ 5.2 Identification of Active and Inactive Constraints
We will now show that the procedure (19) identifies the active constraints. This proof is a straightforward application of Lemma A.2, as explained in what follows. Following a similar reasoning as in [Duchi and Ruan, 2021, Theorem 3], we start by rewriting iteration (19) as
@@ 5.3 Asymptotic Normality
Assumption 5.6(i) is a standard second-order regularity assumption needed to prove the asymptotic normality of SA algorithms Polyak and Juditsky [1992]. In our case, since we are dealing with a joint SA procedure, Assumption 5.6(ii) is a natural symmetric requirement for the $\mu_{n}$ iterates. Assumption 5.6(iii) is a standard restricted strong convexity assumption needed to prove the asymptotic normality of both the SAA and SA procedures Shapiro [1989]; Duchi and Ruan [2021].
@@
Assumption 5.7 is a common assumption needed to ensure that the noise sequence $\{\xi_{k}\}_{k\in\mathbb{N}}$ satisfies the CLT $\frac{1}{\sqrt{n}}\sum_{k=0}^{n-1}\xi_{k}\overset{d}{\to}\mathcal{N}(0,\Sigma)$, for some $\Sigma\geq 0$. In particular, Assumption 5.7(i) is the standard conditional Lindeberg condition needed to prove martingale CLTs (see [Hall and Heyde, 2014, Chapter 3]). Moreover, Assumption 5.7(ii) guarantees that $\Sigma$ is a block-diagonal matrix with $\text{Var}_{X^{(\mu^{\star})}\sim\mathbb{P}_{\mu^{\star}}}\left[G_{\mu^{\star}}(\theta^{\star},X^{(\mu^{\star})})\right]$ and $\text{Var}_{X\sim\mathbb{P}}\left[H(\theta^{\star},\mu^{\star},X)\right]$ on the diagonal.
@@ Acknowledgments
Liviu Aolaritei acknowledges support from the Swiss National Science Foundation through the Postdoc.Mobility Fellowship (grant agreement P500PT_222215). Henry Lam acknowledges support from the InnoHK initiative of the Innovation and Technology Commission of the Hong Kong Special Administrative Region Government, and Laboratory for AI-Powered Financial Technologies. Michael I. Jordan was funded in part by the European Union (ERC-2022-SYG-OCEAN-101071601). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.
@@ Appendix A Technical Preliminaries
Lemma A.1 (Robbins and Siegmund [1971, Theorem 1]).
@@
Lemma A.2 (Duchi and Ruan [2021], Lemma 4.2).
@@
Lemma A.3 (Dembo [2016], Exercise 5.3.35).
@@
Lemma A.4 (Hall and Heyde [2014], Corollary 3.1).
@@ A.1 A Generic Asymptotic Normality Result
To keep the paper self-contained, in what follows we recall the generic asymptotic normality result from Section 13 (pages 12-13) in the Supplementary material of Duchi and Ruan [2021]. There, Duchi and Ruan generalize Polyak and Juditsky’s results Polyak and Juditsky [1992] on asymptotic normality in averaged stochastic gradient methods restricted to an arbitrary subspace of $\mathbb{R}^{s}$. Before stating the result, we need to introduce some notation, as well as two assumptions.
@@
We would like to highlight that Assumption A.5(ii) is slightly weaker compared to Assumption S.A in Section 13 in the Supplementary material of Duchi and Ruan [2021]. Under Assumption 4.1(iii), this weaker version will be enough for us to prove the CLT in Theorem 5.9.
@@
Lemma A.7 (Duchi and Ruan [2021], Proposition S.1).
@@
We highlight the difference between the asymptotic variance in (47) and the asymptotic variance in [Duchi and Ruan, 2021, Proposition S.1 in Supplementary material], which is $(\text{P}H\text{P})^{\dagger}\text{P}\Sigma\text{P}(\text{P}H\text{P})^{\dagger}$. The two expressions are equal. This follows from the fact that P is an orthogonal projection matrix, which guarantees that $\text{P}(\text{P}H\text{P})^{\dagger}=(\text{P}H\text{P})^{\dagger}=(\text{P}H\text{P})^{\dagger}\text{P}$.
@@
The proof follows along similar lines as Duchi and Ruan [2021, Theorems 2 and 3]. For completeness, and since these results are fundamental for proving Theorem 5.2, we provide a full proof for Lemma B.1.
@@
with $\delta_{\Theta}$ the indicator function of the set $\Theta$. Defining $\ell(\theta):=\frac{1}{2}\|\theta\|^{2}+\delta_{\Theta}(\theta)$, we have that the above maximum is precisely the convex conjugate $\ell^{*}(-\sum_{k=0}^{n}\alpha_{k+1}G_{k})$. Now, since $\min_{\theta\in\mathbb{R}^{s}}\{\langle\sum_{k=0}^{n}\alpha_{k+1}G_{k},\theta\rangle+\frac{1}{2}\|\theta\|^{2}+\delta_{\Theta}(\theta)\}$ is the Moreau envelope of $\sum_{k=0}^{n}\alpha_{k+1}G_{k}+\delta_{\Theta}(\theta)$ evaluated at zero, we have that the gradient $\nabla\ell^{*}(-\sum_{k=0}^{n}\alpha_{k+1}G_{k})$ is equal to proximal map evaluated at zero [Rockafellar and Wets, 2009, Theorem 2.26], i.e.,
@@
The proof follows from [Hall and Heyde, 2014, Corollary 3.1], using Assumption 5.7 and the Cramér-Wold theorem. Fix $t\in\mathbb{R}^{s+m}$. Then, recalling the shorthand notation $\xi_{k}$ for the noise vector, we have that